OpenStudy (anonymous):

For graphing- The line through (-3, 4) and (2, -3).

OpenStudy (marissalovescats):

Same thing we did before! Find the slope put that and a point into y-y1=m(x-x1) Then arrange it into y=mx+B and you're done!

OpenStudy (anonymous):

You rock for helping me

OpenStudy (marissalovescats):

It's my pleasure!

OpenStudy (anonymous):

x + y = 6 2x - y = 3 The system of equations a) has no solution b) has one solution c) is coincident

OpenStudy (marissalovescats):

So for this the first thing you want to do is either cancel out your x or y so you can solve for the other. So we'll find x first because as you can see when you combine these two equations the y's will cancel because +y-y=0